The controller of Harrington Company estimates sales and production for the first four months of 2020 as follows: Sales ProduWhat is the change in the cash balance for the period January-March? Cash balance change $
Answer #1

Answer:

Total cash sales $         50840
Accounts receivables balance $         12816
Direct materials inventory balance                 4320 kg
Materials purchases for February $         30500
Cash payments on account $         30440
Ending balance in accounts payable $         27360
Cash balance change $         32,906

Calculations and explanations:

Total cash sales = 0.4*(32700+41000+53400) = 50840
Accounts receivable balance = (0.6*53400)*0.4 =12816

Direct materials inventory balance = 2700*4*0.4 = 4320

Material purchases for February:

Jan Feb March Total
Units to be produced                1100          1400          2000            4500
Kg per unit                     4                  4                  4                    4
Total required for production            4400          5600          8000 18000
Add: desired ending inventory            2700          3200          4320            4320
Total raw materials required            7100          8800       12320          22320
less: beginning inventory            1020          2700          3200            1020
Direct materials purchase            6080          6100          9120          21300
Cost of kg                     5                  5                  5                    5
Total cost of materials purchase          30400       30500       45600       106500

Cash payments on account = (0.6*30400)+(0.4*30500) = 30440

Ending balance in accounts payable = 0.6*45600= 27360

Cash balance change = 50840+ [(0.6*(32700+41000+53400) - 12816] - (106500- 27360)

=35,144